### v1

# SEA: Self-Evolution Agent with Step-wise Reward for Computer Use

Abstract
Computer use agent is an emerging area in artificial intelligence that aims to operate the computers to achieve the user’s tasks, which attracts a lot of attention from both industry and academia. However, the present agents’ performance is far from being used. In this paper, we propose the Self-Evolution Agent (SEA) for computer use, and to develop this agent, we propose creative methods in data generation, reinforcement learning, and model enhancement. Specifically, we first propose an automatic pipeline to generate the verifiable trajectory for training. And then, we propose efficient step-wise reinforcement learning to alleviate the significant computational requirements for long-horizon training. In the end, we propose the enhancement method to merge the grounding and planning ability into one model without any extra training. Accordingly, based on our proposed innovation of data generation, training strategy, and enhancement, we get the Self-evolution Agent (SEA) for computer use with only 7B parameters, which outperforms models with the same number of parameters and has comparable performance to larger ones. We will make the models’ weight and related codes open-source in the future.
footnotetext: Corresponding author: Yukang Huo. Email: yukanghuo.ai@gmail.com
I Introduction
Computer use agent has been an attractive technology that follows the users’ instructions to manipulate computers and match the users’ expectations, and it has been a milestone towards general artificial intelligence. In order to achieve the goal of computer use, the agent should understand the computer environment, provide a plan based on the instructions, and finally manipulate the computer. For the computer environment, researchers will take a screenshot as a representation of that. With the development of the MLLMs [1, 2], they have the remarkable ability of perception and reasoning, which can meet the requirements of a computer-use agent.
Figure 1: The comparison with the SOTA methods
Despite significant progress in the field of intelligent agents, building robust agents for real-world deployment remains highly challenging, particularly in scenarios that involve complex multimodal reasoning and dynamic, uncertain environments. One of the most critical challenges in training intelligent agents is acquiring high-quality data. Current mainstream data construction approaches predominantly rely on manually designed task templates and annotations. This process is not only costly and difficult to scale but also often fails to meet the semantic coverage and diversity required by complex tasks. Moreover, real-world tasks typically involve multi-step, long-horizon interactions, where informative reward signals only emerge after completing the entire task or passing through several key steps. Such sparse reward settings make it difficult for the agent to establish a clear correlation between individual actions and final outcomes, leading to suboptimal task performance. However, when handling complex GUI tasks, encoding image data often requires processing a large sequence of tokens, many of which contribute minimally to task execution and reasoning but incur significant computational costs. In addition, existing reasoning mechanisms frequently depend on external models to assist in evaluating and selecting trajectories. This reliance introduces significant computational overhead and increases system complexity, further limiting the efficiency and scalability of the overall agent architecture.
To address these challenges, we propose a set of novel agent training methods, designed to enable efficient and scalable task learning and execution. In the data acquisition phase, we propose a closed-loop pipeline to generate the verifiable tasks, which leverages a task-generation agent and a code-generation agent to automatically synthesize tasks along with their corresponding execution and verification Python programs. This dual-code generation strategy ensures that each task is accompanied by a well-defined success criterion and a programmatic validation method, effectively forming a closed-loop supervision structure. By tightly coupling the task definition with its success condition, the proposed pipeline significantly enhances both the effectiveness and usability of the generated data. Furthermore, we use virtual environment to run each task using its corresponding execution program and applying the verification program to retain only those tasks that are successfully completed. And then we propose the Generation and Assessment for Trajectory Extraction (GATE) method, designed to construct high-quality and representative training trajectories that improve agent learning efficiency. Specifically, for each validated task, we perform multiple rounds of inference to generate diverse candidate trajectories. As training progresses, we gradually replace the inference model with our agent during the middle and later stages of training. Each generated trajectory is then filtered again using the verification code to ensure successful task completion. Among the successful trajectories for the same task, we prioritize those with the fewest required steps, selecting more efficient execution paths. Additionally, we train a step filtering model to eliminate redundant or incorrect steps within each trajectory, resulting in high-quality, concise training data.
During the training phase, we introduce the Trajectory Reasoning by Step-wise Reinforcement Learning(TR-SRL). This component evaluates each step in a task trajectory by comparing the agent’s execution results against the corresponding ground-truth outcomes. In addition, TR-SRL assesses the consistency between the agent’s generated thoughts and executed actions at each step, ensuring alignment between internal reasoning and external behavior.
At last, we propose the method to further enhance the generalization ability and effect of the model. Specifically, we first train a grounding model and integrate its weights, which effectively improves the performance of both grounding and planning. In addition, we introduce a Temporal Compressed Sensing Mechanism to optimize perception efficiency by modeling the importance of input image tokens over time. By retaining only the most important tokens, the model achieves a significant reduction in training cost while maintaining critical semantic content and improving inference efficiency.
Our main contributions can be summarized as follows:
  • We propose a data generation pipeline to produce the closed-loop verifiable tasks’ trajectory for reinforcement learning. The proposed pipeline will simultaneously generate the task instruction, execution program, and verification program. To further refine data quality, we introduce the GATE, which identifies optimal training steps via trajectory sampling and evaluation.
  • We propose the TR-SRL(Trajectory Reasoning by Step-wise Reinforcement Learning), which uses the step-wise training manner to replace the long-horizon training of manipulating the computer.
  • We proposed a Grounding-Based Generalization Enhancement method to combine the planning ability and grounding ability into a unified model. The proposed method not only can keep the original models’ performance but also can improve it.
In summary, based on these proposed methods including verifiable data generation, efficient reinforcement learning, and model enhancement, we create the SEA, which is the Self-Evolution Agent for computer use. As shown in the 1, our agent outperforms the 7B models and has a comparable performance with a larger number of parameters.
II Related Works
II-A High-quality Training Data
High-quality data is essential for bridging the gap between the symbolic world [3] and the digital world [4], thereby advancing the development of intelligent agents. Early works such as Rico [5] and Mini-Wob [6] provided serialized GUI data and low-level keyboard/mouse interactions for mobile applications and web-based tasks, respectively. Subsequent studies have further expanded data resources for mobile [7, 8, 9, 10], web [11, 12, 13], and desktop applications [14]. To train agents more efficiently, researchers typically use trajectory data, which consists of sequences containing GUI information, including both low-level and high-level instructions and corresponding operations [15, 16, 17]. In addition, some studies utilize demonstration data, filtered imitation learning, and self-collected data [18, 19] to fine-tune vision-language models (VLMs), such as AutoUI and CogAgent [20, 21]. PAE [22] generates tasks automatically to provide diverse training data for agents.
However, these methods face several challenges in practical applications: the reliance on costly manually annotated task templates, the inability of static data to adapt to real-world dynamics, and the inconsistent quality of task trajectories—all of which severely limit the effectiveness and adaptability of agents in diverse environments. So in our paper, we have proposed a pipeline to generate high-quality verifiable tasks with correct trajectory.
II-B Compute Use Agent
Recent advancements in multi-modal large language models (MLLMs) have enabled agents to perform increasingly complex tasks within graphical user interfaces (GUIs). Models such as SeeClick [23] and ScreenAgent [24] leverage vision-language models (VLMs) to interpret UI screens and execute interactive tasks based on multimodal prompts. GUI-R1 [25] enhances UI grounding performance by introducing reinforcement learning mechanisms. Song [26] proposes a framework for API-driven web agents that utilizes task-specific background knowledge to execute complex web operations. The AgentWorkflow Memory (AWM) module [27] further optimizes memory management by selectively providing relevant workflows to guide the agent’s subsequent actions. Meanwhile, there are some specialized computer-use agents, such as PresentAgent [28], that aim to help users generate the presentation. Despite the substantial progress made in perception, understanding, and execution, existing approaches still face key limitations: sparse rewards in long-horizon tasks, high computational costs during training, and a lack of modeling for the consistency between thought and action within action sequences. So in this paper, we propose to use the step-wise RL training to replace the long-horizon training, which is achieved by the dedicated designing of rewards and the advanced RL training strategy.
II-C Reinforcement Learning
As the fascinating performance of the DeepSeek-R1 [29], researchers have noticed that the great of the Post-training in improving the performance of the LLM, especially comparing with the minor imporvence of the pre-trainig phase. OpenAI’s o1 [30] and DeepSeek-R1 [29] demonstrate strong capabilities in tasks such as mathematical reasoning [31], code generation [32], and multimodal reasoning [33, 34] through structured reward signals. Sweet-RL [35] introduces a multi-round DPO framework to improve long-horizon behavior in language agents, while ARPO [36] increases success rates on complex multi-turn GUI tasks by incorporating experience replay and task selection strategies. DISTRL [37] boosts training efficiency and success rates for GUI agents in mobile device control tasks through its novel asynchronous distributed reinforcement learning framework and the custom A-RIDE algorithm, while also enhancing generalization across diverse tasks. UI-TARS [38] significantly improves performance in complex GUI tasks through innovations in enhanced perception, unified action modeling, and dual-stage reasoning. Moreover, some methods propose interactive refinement of LLMs/VLMs—particularly Web/GUI agents—using self-evaluated feedback [18, 39, 40, 37].
III Method
In this section, we provide a detailed description of the proposed method. Firstly, we will provide a simple review of reinforcement learning. We will then dive into our proposed method. As we have mentioned, the proposed method consists of three parts: automatic data collection, RL with multiple rewards, and model enhancement with a grounding model.
III-A Data Engine for Self-evolution
One of the key points to make a model evolve is that we could have the ability to scale up the training data, which will provide the model more probability to explore. Nevertheless, the collection of this kind of data, operating the computer in a human-style, is so complex that the amount of this data is insufficient for scaling up. To solve this challenge, we propose a novel method to generate a trajectory dataset of manipulating different computer software. Specifically, the proposed method can be divided into three parts: (1)Generation of closed-loop verifiable tasks; (2) Generation and Assessment for Trajectory Extract; (3) Filtering of data with a step model.
III-A1 Generation of closed-loop verifiable tasks
To enable the agent to better handle complex multimodal tasks and dynamic environment interactions, we design a closed-loop task data generation pipeline, as illustrated in Figure 2. Firstly, the proposed pipeline will use a Task Agent to produce many task instructions. The task agent will use some actual tasks of a specific software as the few-shot examples to evoke the agent’s in-context learning of the target sofware. During the generation process, the task agent will periodically check whether the generated tasks can be executed and verified and whether there are some similar tasks have been generated. Meanwhile, the agent will produce some guidelines of how to complete the task. Secondly, the code generation agent will take the generated task instruction and the guideline as inputs to generate a batch of execution and verification program. For execution program, it’s target is to achieve tasks’ goals described in instructions by using the Python. And for the verification program, it judges whether the execution programs’ output matches the tasks’ requirement. In the end, we will automatically run pairs of execution program and verification to select the correct verification program. So in this way, we can get the verifiable tasks.
Figure 2: Illustration of Generating Closed-loop Task Data.
Figure 3: Illustration of the data generation and multi-stage trajectory filtering strategy.
III-A2 Generation and Assessment for Trajectory Extraction
To enable the agent to learn the most relevant and effective action sequences, we propose a novel multi-stage trajectory filtering strategy, as illustrated in Figure 3. After getting the task instruction and verification program, we will firstly use our cold-start model, UI-TARS-1.5-7B [38], to perform multiple rounds of inference and obtain diverse task trajectories. And the verification program is then used to select only those trajectories that lead to successful task completion. For each task, we prioritize trajectories with the minimal number of steps based on step-length information. Additionally, to further ensure the quality of each step within the trajectory, we train a step filtering model to filter out redundant or erroneous steps, thereby yielding optimal training trajectories. And the step model will be introduced in the next section. It is notable that during the training process, we will gradually replace the cold-start model with our trained model to generate trajectories for later training stages. In this way, we can make our model evolve through the whole training phase.
III-B Filtering data with step model
For the step filtering model, its task is to determine whether the action has been successfully executed based on the two screenshots before and after this step, the thought and action information. We manually annotated some ground truth data and continuously adjusted prompts to ensure that the model’s performance follows our filtering criteria. Then, we randomly extracted 10,000 step data from the task trajectory and used the QWen2.5-VL-72B [1] model and adjusted the prompt for inference. Each piece of data was subjected to multiple inferences and voting to ensure accuracy as much as possible, and finally used to train the step filtering model. Through this method, we distill the capabilities of the large model into the small model, making the step filtering model lighter while maintaining excellent performance consistent with the large model.
Figure 4: Illustration of Trajectory Reasoning by Step-wise Reinforcement Learning.
III-C Step-wise Reinforcement Learning
In this part, we will introduce Trajectory Reasoning by Step-wise Reinforcement Learning. Specifically, we will use the UI-TARS-7B [38] as our cold-start model, and based on this model, we will use the proposed methods to produce the final agent.
III-C1 Preliminary of GRPO
Although there are lots of improvements in the GRPO, in this paper, we still choose it, as a result of its balance between performance and tidiness. Group Relative Policy Optimization (GRPO) [31] is a reinforcement learning method designed to the efficient fine-tuning of large language models (LLMs) without relying on an explicit value function or critic. GRPO replaces the need for a trained value function in standard Proximal Policy Optimization (PPO) with the average rewards of multiple outputs for the same problem, making it particularly suitable for LLMs.
Given a batch of $G$ responses $\{o_{i}\}_{i=1}^{G}$ from a query $q$, each consisting of a sequence of tokens $o_{i}=(o_{i}(1),\dots,o_{i}(T))$, the GRPO objective is defined as:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}[\begin{subarray}{c}q\sim\mathcal{P}(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)\end{subarray}]$ (1)
$\quad\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Bigg{\{}\min\Bigg{[}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i,t},$
$\qquad\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}\Bigg{]}$
$\qquad-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]\Bigg{\}}$
Where $\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]$ is shown in the following equation 2. $\hat{A}_{i,t}$ is the group-normalized advantage for token t in response $o_{i}$, computed as equation 3.
$\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]=\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-\log\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-1$ (2)
$\hat{A}_{i,t}=\frac{r_{i}-\mu}{\sigma},\quad\text{with }r_{i}\text{ the total reward of }o_{i}$ (3)
Where $\mu$ and $\sigma$ are the mean and standard deviation of rewards in the group.
III-C2 Step-wise RL training
Unlike single-step reinforcement learning, computer use agents are required to perform multi-step reasoning and decision-making while interacting with dynamic environments that provide visual feedback. However, training on long-horizon tasks often suffers from the problem of sparse rewards. To address this issue, we propose a step-wise training method, as illustrated in Figure 4. In our training setup, each task trajectory consists of multiple steps, where each step comprises a screen observation $s_{i}$, a keyboard or mouse action $a_{i}$, and an associated reward $r_{i}$, with $i$ denoting any step in the trajectory.
III-C3 Rewards
To predict the next action $a_{i+1}$, the model takes as input the user’s request (i.e., task goal), the current screen observation, and the historical sequence of actions and states. As shown in the  4, we have used multiple rewards, specifically, we introduce three types of reward mechanisms to optimize the training process:
Step Reward. For each step in a trajectory, we define a step-level reward based on whether it succeeds. If the agent completes a step and the execution result matches the ground truth, the reward is set to $r_{i}=1$; otherwise, $r_{i}=0$. This step-wise reward structure is particularly beneficial for long-horizon tasks, as it allows the agent to learn effective action sequences over multiple interactions. By providing clear and immediate feedback at each step, the agent can better identify which actions are effective and which require adjustment.
Reasoning and Action Consistency Reward. During training, every action produced by the agent is accompanied by a corresponding thought process. If the predicted action is consistent with the reasoning (i.e., aligned in logic and intent), a consistency reward $r_{c}=1$ is given; otherwise, $r_{c}=0$. This mechanism effectively constrains the model to maintain alignment between its reasoning and the resulting action, leading to more coherent and reliable behavior.
Action Format Reward. During each rollout, responses from the agent are parsed into discrete actions. If a response fails to conform to the required action schema and cannot be parsed, we assign a penalty of $r_{f}=0$, otherwise $r_{f}=1$. This encourages the model to generate syntactically valid and executable actions.
III-C4 Training Objective
We divide each task trajectory $\tau=(s_{0},a_{0},\ldots,s_{n},a_{n})$ into multiple steps. At each step, the agent observes the screen state $s_{i}$ and generates an action $a_{i}$ to fulfill the task instruction $x_{i}\in D$. Our training objective is to maximize the expected reward for step completion, consistency between thoughts and actions, and action format:
$\max_{\theta}\mathbb{E}_{x_{i}\sim D,\tau_{i}\sim\pi_{\theta}}[r_{i}(x_{i},\tau_{i})+r_{c}(x_{i},\tau_{i})+r_{f}(x_{i},\tau_{i})]$ (4)
We optimize this objective using GRPO, which estimates token-level advantages via group-normalized trajectory rewards, allowing efficient and scalable training without a value function.
III-D Inference Stage
During the inference stage, we build upon the GTA1 [41] and replace the original inference model with our proposed model SEA. As a result, both the sampling and selection of candidate trajectories during inference are handled entirely by SEA, without relying on any external scoring models. This design significantly reduces computational overhead and effectively improves decision-making efficiency.
TABLE I: Comparison of Various Planners and Grounding Methods on ScreenSpot-Pro.
Agent Model Development Creative CAD Scientific Office OS Avg
Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg
QwenVL-7B [42] 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.7 0.0 0.4 0.0 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.1
GPT-4o [43] 1.3 0.0 0.7 1.0 0.0 0.6 2.0 0.0 1.5 2.1 0.0 1.2 1.1 0.0 0.9 0.0 0.0 0.0 1.3 0.0 0.8
SeeClick [23] 0.6 0.0 0.3 1.0 0.0 0.6 2.5 0.0 1.9 3.5 0.0 2.0 1.1 0.0 0.9 2.5 0.0 1.5 1.8 0.0 1.1
Qwen2-VL-7B [44] 2.6 0.0 1.3 1.5 0.0 0.9 0.5 0.0 0.4 6.3 0.0 3.5 3.4 1.9 3.0 0.9 0.0 0.5 2.5 0.2 1.6
OS-Atlas-4B [45] 7.1 0.0 3.7 3.0 1.4 2.3 2.0 0.0 1.5 9.0 5.5 7.5 5.1 3.8 4.8 5.6 0.0 3.1 5.0 1.7 3.7
ShowUI-2B [46] 16.9 1.4 9.4 9.1 0.0 5.3 2.5 0.0 1.9 13.2 7.3 10.6 15.3 7.5 13.5 10.3 2.2 6.6 10.8 2.6 7.7
CogAgent-18B [47] 14.9 0.7 8.0 9.6 0.0 5.6 7.1 3.1 6.1 22.2 1.8 13.4 13.0 0.0 10.0 5.6 0.0 3.1 12.0 0.8 7.7
Aria-UI [48] 16.2 0.0 8.4 23.7 2.1 14.7 7.6 1.6 6.1 27.1 6.4 18.1 20.3 1.9 16.1 4.7 0.0 2.6 17.1 2.0 11.3
UGround-7B [49] 26.6 2.1 14.7 27.3 2.8 17.0 14.2 1.6 11.1 31.9 2.7 19.3 31.6 11.3 27.0 17.8 0.0 9.7 25.0 2.8 16.5
Claude Computer Use 22.0 3.9 12.6 25.9 3.4 16.8 14.5 3.7 11.9 33.9 15.8 25.8 30.1 16.3 26.9 11.0 4.5 8.1 23.4 7.1 17.1
OS-Atlas-7B [45] 33.1 1.4 17.7 28.8 2.8 17.9 12.2 4.7 10.3 37.5 7.3 24.4 33.9 5.7 27.4 27.1 4.5 16.8 28.1 4.0 18.9
UGround-V1-7B [49] 35.5 27.8 13.5 38.8 48.8 26.1 31.1
UI-TARS-2B [38] 47.4 4.1 26.4 42.9 6.3 27.6 17.8 4.7 14.6 56.9 17.3 39.8 50.3 17.0 42.6 21.5 5.6 14.3 39.6 8.4 27.7
UI-TARS-7B [38] 58.4 12.4 36.1 50.0 9.1 32.8 20.8 9.4 18.0 63.9 31.8 50.0 63.3 20.8 53.5 30.8 16.9 24.5 47.8 16.2 35.7
UI-TARS-72B [38] 63.0 17.3 40.8 57.1 15.4 39.6 18.8 12.5 17.2 64.6 20.9 45.7 63.3 26.4 54.8 42.1 15.7 30.1 50.9 17.5 38.1
Ours(Grounding Model) 76.0 21.4 49.5 61.6 19.6 44.0 45.2 18.8 38.7 80.6 31.8 59.4 84.2 54.7 77.4 57.9 33.7 46.9 67.0 27.3 51.9
III-E Grounding-Based Generalization Enhancement.
After using the step-wise RL method to train the cold-start model using the verifiable tasks collected through the proposed data engine, we can get a model with strong planning ability. However, the computer use agent not only needs the planning also needs the grounding to locate target elements in the screen. In order to enhance the model’s perception ability, we firstly train a grounding model and then merge the planning model with the grounding model. In the end, we use the Temporal Compressed Sensing Mechanism to make agent pay more attention of images.
III-E1 Train Grounding Model
We will use the supervised fine-tuning to get a grounding model, and the training data is one of the key points influencing its performance. At present, there are numerous open-source datasets in the field of grounding.
Aguvis [50] and OS-Atlas[45] have organized and filtered many previous open-source datasets with high quality, so we have conducted our data filtering based on them. We select 30000 pieces of data from each subset of these two datasets (if fewer than 30000 pieces, we take all), and these 30000 pieces of data should come from different screenshots, rather than other elements of the same screenshot. We used these different datasets and the same configuration for SFT training, and evaluated them using screenspot-v2. Finally, we selected five higher-quality subsets from two large datasets as the training set for the grounding model. Through these experiments, we also found that the quality of the dataset is far more critical than the quantity.
In addition, detailed instructions are helpful for the grounding task. To reconstruct the data samples, we retain the first two steps from [51] For sample construction, we retain the first two steps from [51] while optimizing the third step. The original method discards samples with imprecise predicted coordinates during sample construction. However, we observe that these discarded samples often correspond to more complex tasks, involving richer operational logic and reasoning paths. Therefore, we preserve the original task instructions (the original commands) of these samples and reincorporate them into the training set. To further enhance data diversity and model robustness, we apply data augmentation techniques such as image scaling and color inversion to the filtered dataset.
Finally, we train our grounding model using SFT and GRPO reinforcement learning, employing the following IoU-based loss or reward function:
$\text{Reward}=\min\Bigg{(}1,$ $\text{IoU}(B_{p},B_{gt})\cdot$ (5)
$\frac{1}{\mathbb{I}\left[\text{IoU}(B_{p},B_{gt})\leq 0.7\right]+\varepsilon}\Bigg{)}$
Here, $\mathbb{I}$ denotes the indicator function, which equals 1 if the condition inside the parentheses is satisfied, and 0 otherwise. $\text{IoU}(B_{p},B_{gt})=\frac{|B_{p}\cap B_{gt}|}{|B_{p}\cup B_{gt}|}$ represents the Intersection over Union (IoU) score between the predicted bounding box $B_{p}$ and the ground truth box $B_{gt}$. $\varepsilon$ is a small constant (e.g., $10^{-6}$) used to prevent division by zero.
III-E2 Model Merge
After getting the grounding model, we need to find a simple and effective way to combine the planning and grounding abilities into one model. Meanwhile, we do not want to use extra computation and data to distill these two models into one. As a result of that, we leverage the DARE [52] to merge these two models. Based on the idea of task vectors, this method sets most incremental parameters (i.e., the differences between parameters of the grounding model and planning model) to zero without affecting the performance of original models. DARE randomly discards incremental parameters at a rate of $p$ and rescales the remaining parameters by 1/(1p)1/(1-p) to approximate the original embedding.
III-F Temporal Compressed Sensing Mechanism
To improve training efficiency and reduce computational cost, we propose the TCSM (Temporal Compressed Sensing Mechanism). Specifically, TCSM first differentiates tokens based on their recency. Newer tokens typically carry more information about the current task state and objectives, and thus are assigned higher priority during training. Tokens with relatively lower priority undergo structured compression to reduce computational overhead.
$\mathcal{I}_{\text{TCSM}}=\left\{\operatorname{Pad}\left(\operatorname{Resize}(I_{i})\right)\ \middle|\ i\in\mathcal{H},\ i\geq n-k+1\right\}$ (6)
Here, $\mathcal{H}$ denotes the sequence of historical images, containing $n$ frames in total. $I_{i}$ represents the $i$-th frame in the history. To focus on the key temporal context, we retain only the most recent $k$ frames, while the remaining earlier frames are compressed for efficient representation.
IV Experiments
In this section, extensive experiments are conducted to compare the proposed agent with other state-of-the-art methods, followed by some relevant analysis of our method.
IV-A Experimental Setup
Implementation Details. For the training part, we use the LLaMa-Factory to train our grounding model which is used in the model integration. And for the whole model, we use the Easy-R1 to train the model by using the GRPO. During the whole training phase, we use one server with 8xA100(80GB) GPU. During the inference phase, we reference the GTA-1 [41] to have better results. Firstly, we used our agent to get $N$ possible results, in our experiment, $N$ equals to 8. And then we will use another model to select the best one from them, which is different from GTA-1 [41] using the GPT-o3.
Datasets and Metrics. As we have mentioned before, we will use the generated verifiable tasks to produce the trajectory training data. In our experiments, we use the 400 tasks. And for each inference in the trajectory, we produce eight results. For the evaluation, we evaluate our improved grounding model on the ScreenSpot-V2 [23, 45] and ScreenSpotPro [53] datasets using accuracy as the evaluation metric. Subsequently, we evaluate our proposed model on the OSWorld benchmark [54], which is designed to evaluate the computer use agent’s performance on open-ended GUI tasks. This benchmark comprises 369 tasks spanning real-world web and desktop applications, providing a diverse and challenging platform to evaluate the agent’s ability to complete user tasks within a Linux environment. Performance is measured primarily by task success rate.
IV-B Experimental Results
IV-B1 Grounding Performance
We first evaluate the performance of our improved and trained grounding model. Among models of comparable scale, our grounding model achieves SOTA performance, as shown in  I.
IV-B2 OS-World Performance
We compare our method with state-of-the-art approaches on the OSWorld benchmark, with results summarized in II. As shown, despite utilizing significantly fewer parameters and incurring lower training costs, the proposed method still surpasses the larger model UI-TARS-72B-DPO [38] by 3.0% in task success rate. Furthermore, we reproduced the baseline model UI-TARS-1.5-7B [55] under the same experimental conditions. Compared to this baseline, SEA achieved a performance improvement of 5.2% and attained the highest task success rate among models of the same scale on the OSWorld benchmark, outperforming all existing state-of-the-art approaches. These results demonstrate the effectiveness of the proposed method in handling complex and realistic user tasks across diverse scenarios, highlighting the robustness and generalization capability of our approach. Meanwhile, as shown in this table, the model without the enhancement also achieves the SOTA performance, which proves the effectiveness of the data generation of verifiable tasks and the step-wise reinforcement learning.
TABLE II: Performance Comparison of the Proposed Method and other SOTA Models on the OSWorld Benchmark.
Model OSWorld (%)
OpenCUA-3B [56] 18.8
OpenCUA-Qwen2-7B [56] 23.1
OpenCUA-7B [56] 26.6
UI-TARS-72B-DPO [38] 27.1
UI-TARS-1.5-7B [55] 27.4 (24.9)
Ours (w/o enhancement) 28.1
Ours 30.1
IV-C Ablation Study
Generalization Enhancement on planning. As we have mentioned, we propose the enhancement method to improve the generalization of the model trained by RL. As presented in  II, the task completion rate increases from 28.1% to 30.1%, representing an improvement of 2.0%. This indicates that by integrating weights from different task modules and adopting a self-evolution mechanism, the model can achieve better generalization to novel task scenarios. We think that this kind of enhancement makes the original model have more knowledge of the PC’s screenshots, which will help the model produce more precise planning of different software, even if it is a new application.
V Conclusion
In this paper, we propose a series of innovations to produce a self-evolution for automatically operating the computers following the user’s instructions. We have proposed a novel pipeline to generate closed-loop verifiable tasks’ trajectories for training agents’ ability in reasoning, which solves the challenge of letting the model evolve itself from the data. In this part, we use two agents to generate the task instruction and the verification program, respectively. And then the generated verifiable tasks will be passed to extract the trajectory with assessment, which periodically replaces the generation model with the latest trained model and produces high-quality data. After this, the filter model will select the data for training. Secondly, we have proposed a step-wise reinforcement learning method to replace the long-horizon training, which frees the model to evolve from the computation limitation. Thirdly, we have proposed a enhancement method to help the model have the great grounding and planning ability at the time.
Experiments on the OSWorld benchmark demonstrate that the proposed agent achieves competitive results, which outperform other agents of the same scale and have comparable results to larger models. And we hope that our proposed method, including the pipeline to generate verifiable task trajectories, trajectory reasoning by step- wise reinforcement learning and grounding-based generalization enhancement can offer a promising new direction for the computer use agent area.